### v1

# Audio–Visual Speech Enhancement: Architectural Design and Deployment Strategies

Abstract
This paper introduce a new AI-based Audio-Visual Speech Enhancement (AVSE) and presents a comparative performance analysis of deployment architectures. The proposed AVSE system integrates convolutional neural networks (CNNs) for spectral feature extraction and long short-term memory (LSTM) networks for temporal modeling, enabling robust speech enhancement through multimodal fusion of audio and visual cues. We investigate multiple deployment scenarios, including cloud-based, edge-assisted, and standalone device implementations, evaluating their performance in terms of speech quality improvement, latency, and computational overhead. Real-world experiments are conducted across heterogeneous network conditions (Ethernet, Wi-Fi, 4G, and 5G) to assess the trade-offs between processing delay, communication latency, and perceptual speech quality. Results show that while cloud deployment achieves superior enhancement quality, edge-assisted architectures offer the best compromise between latency and intelligibility, meeting real-time requirements under 5G and Wi-Fi 6 conditions. The findings provide practical guidelines for selecting and optimizing AVSE deployment architectures in diverse application domains, from assistive hearing devices to telepresence and industrial communications.
Index Terms:
Audio–Visual Speech Enhancement, Deep Learning, Deployment Architecture, Comparative Performance, COG-MHEAR, Edge Computing, Cloud Processing.
I INTRODUCTION
Audio-Visual Speech Enhancement (AVSE) leverages the complementary nature of auditory and visual cues to improve speech perception in adverse acoustic conditions. In humans, this multimodal integration is an innate cognitive process, enabling effective communication in the presence of background noise, competing talkers, or reverberant environments. Neuroscientific studies have established the value of visual input in selective auditory attention. Golumbic et al. [13] demonstrated that visual access to a talker’s face significantly improves the auditory cortex’s ability to segregate and track the target speaker in multi-speaker environments, a phenomenon often referred to as the “Cocktail Party” effect. Likewise, McGurk and MacDonald [3] revealed that incongruent lip movements can alter auditory perception entirely, demonstrating that speech understanding is not purely an auditory process but is strongly influenced by visual context. Further, Partan and Marler [9] emphasized that multimodal communication signals, common in group-living species, improve clarity, robustness, and interpretation.
Motivated by these findings, AVSE systems employ visual information—such as lip motion, facial dynamics, and articulatory patterns—to guide audio enhancement. The integration of these modalities allows for more accurate speaker separation, noise suppression, and reverberation reduction. Modern deep learning techniques have accelerated this field by enabling hierarchical feature extraction and cross-modal fusion at multiple abstraction levels. Convolutional Neural Networks (CNNs) are typically employed to extract discriminative visual features, while spectrogram-based CNNs or waveform-level encoders capture audio representations. Recurrent architectures, such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs), are frequently used to model temporal dependencies, which are critical for capturing speech dynamics.
The applications of AVSE extend beyond traditional hearing aids. In military communications, AVSE can enhance intelligibility in complex acoustic environments such as the battlefield. In consumer electronics, it improves voice-command recognition for smart assistants and in-car systems, even in noisy driving conditions [27]. In education and telepresence, AVSE can ensure intelligible communication in virtual classrooms or video conferencing, while in industrial and public-service robotics, it enables reliable interaction despite background noise. The healthcare sector can also benefit, as AVSE improves telemedicine consultations for patients with speech or hearing impairments [26].
While algorithmic design is central to AVSE performance, deployment architecture has a decisive role in real-world effectiveness. Processing location—whether cloud-based, edge-assisted, or entirely on-device—affects latency, speech quality, and system scalability. Communication bandwidth, device computational capability, and network stability determine whether a deployment can meet the real-time requirements for interactive speech enhancement. In particular, deployment trade-offs become critical when serving resource-constrained devices such as IoT endpoints, low-cost smart glasses, or wearable hearing-assist devices.
In this paper, we present a complete AVSE system implemented within the COG-MHEAR framework. The system uses a CNN–LSTM-based fusion architecture, and we evaluate its performance across multiple deployment configurations: server-side cloud processing, edge-assisted processing, and local execution. The contributions of this work are threefold:
  1. 1.
    We propose an AVSE architecture that combines convolutional and recurrent feature extraction with intermediate fusion, enabling robust speech enhancement in challenging conditions.
  2. 2.
    We design and implement a flexible deployment framework capable of supporting heterogeneous device capabilities, from lightweight IoT hardware to high-performance computing platforms.
  3. 3.
    We conduct a comparative performance analysis under realistic network conditions, quantifying trade-offs between latency, intelligibility, and computational load across deployment scenarios.
The remainder of this paper is organized as follows. Section II describes the acoustic signal modeling framework. Section III reviews the visual feature extraction and fusion strategies. Section IV presents the proposed AVSE architecture and deployment designs. Section V reports a comparative evaluation of deployment scenarios, and Section VI concludes with implications for future AVSE systems.
II Acoustic Signal Modeling
Central to the concept of AVSE is the nuanced modeling of acoustic signals. Here, each impulse response emanating from a spatial position of a target source to a microphone is represented as $h_{s}[k]$, where $k$ signifies a discrete-time index. This impulse response is bifurcated into two distinct components: the early part $h_{s}^{1}[k]$, encompassing the direct sound and early reflections, and the late part $h_{s}^{2}[k]$, which pertains to the subsequent reverberations. Consequently, the observed acoustic mixture at any given instance $k$ can be expressed in a mathematical formulation as:
$y[k]=\sum_{s=1}^{T}x_{s}[k]+d[k]$ (1)
$x_{s}[k]=x_{s}^{0}[k]\circledast h_{s}^{1}[k]$ (2)
$d[k]=\sum_{s=1}^{T}x_{s}^{0}[k]\circledast h_{s}^{2}[k]+\sum_{c=1}^{N}d_{c}[k]$ (3)
In this formulation, $x_{s}^{0}[k]$ represents the original speech signal emitted by the $s$-th target speaker. The term $x_{s}[k]$ denotes the clean speech signal from the $s$-th target speaker, inclusive of low-order reflections as captured by the microphone. The component $d_{c}[k]$ symbolizes the noise contribution from various sources as observed at the microphone.
III Visual Information Integration
The incorporation of visual data significantly augments the AVSE process. Studies, such as those mentioned in [25], have demonstrated the critical role visual elements play in enhancing speech recognition systems. Visual features not only provide high-level information about speech patterns but also furnish fine-grained articulatory details. The visual processing initiates with face detection and tracking algorithms [11], [17], [10], [4], [7], which are vital for delineating input dimensionality and discerning the count of speakers. This step significantly reduces the complexity of subsequent stages by focusing solely on the target faces.
III-A Deep Learning in AVSE
The actual fusion and processing of the audio and visual data in AVSE are achieved using an ensemble of deep neural network models. These models, by virtue of their design, can learn and synthesize representations of the acoustic and visual modalities at several levels of abstraction. Convolutional Neural Networks (CNNs) [6], [16], [18], [19], [24], with their inherent characteristics of sparse connectivity and translation invariance, are typically employed to process visual frames and extract pertinent features. They also find applications in acoustic signal processing, handling either the spectrogram or the raw waveform. The training of these networks, often conducted using stochastic gradient descent (SGD) [2], [1], RmsProp [12], and Adam [15], plays a critical role in parameter optimization and model effectiveness.
III-B Fusion Paradigms With Deep Learning
In the realm of AVSE, various strategies, including early and late fusion [23], [21], [22], are employed. Nonetheless, the predominant technique is intermediate fusion [20], where different unimodal representations amalgamate in a single hidden layer. This paper focuses on concatenation fusion, wherein audio and visual feature maps merge at corresponding spatio-temporal locations, allowing subsequent layers to decipher and integrate the information effectively. The fusion function, mathematically represented as $f:y_{l}[k],v_{l}[k^{\prime}]\rightarrow x_{l}[k]$, combines the audio feature map $y_{l}[k]$ and the video feature map $v_{l}[k^{\prime}]$ at times $k$ and $k^{\prime}$, respectively, producing an output map $x_{l}[k]$ at the $l$-th layer of the DNN. $x_{l}^{\text{cat}}[k]=f_{c}at(y_{l}[k],v_{l}[k^{\prime}])$ stacks the two feature maps at the same spatio-temporal locations $i,j,k$ across the feature channels $d$:
${x_{l}^{\text{cat}}[k]}_{i,j,2d}={y_{l}}_{i,j,d}[k]\quad\text{and}\quad{x_{l}^{\text{cat}}[k]}_{i,j,2d-1}={v_{l}}_{i,j,d}[k^{\prime}]$ (4)
where $x_{l}^{\text{cat}}[k]\in\mathbb{R}^{H\times W\times 2D}$.
Concatenation does not inherently define a correspondence but leaves it to subsequent layers to define, by learning suitable filters that weight the layers.
III-C Recurrent Neural Networks in AVSE
Given the inherently sequential nature of speech signals, Recurrent Neural Networks (RNNs) [5], Long Short-Term Memory networks (LSTMs) [5], [8], and Gated Recurrent Units (GRUs) [14] are particularly suitable for AVSE. These networks are adept at processing data where temporal dynamics are significant, making them ideal for speech signals where temporal coherence is paramount.
IV Proposed Solution
Figure 1 presents a pragmatic framework for the deployment of the developed AVSE algorithm within the context of real-world applications. In order to enable the service in lowquality terminals (e.g., IoT devices), the remote server takes over all the processing, including the preprocessing of the raw data. By doing so, any device with a camera, microphone, and communication capabilities can use the service without any requirements on its computing power. This allows us to apply the service to some cheap IoT devices to achieve specific functions. For high-performance devices (e.g., smartphones), offloading all processing to the server can also reduce energy consumption and improve battery life.
Figure 1: A block diagram of enabling real-world COG-MHEAR service on terminal devices.
As the figure depicts, the client terminals only collect the raw audio and video, then upload them to the server and wait for the feedback from the server. It should be noted that collecting and playing audio requires parallel processing, so we use two independent threads. The server will listen to clients and wait for the raw audio-video data from clients. It starts processing using the stored trained model once it listens to data transmission from the client. Considering that processing latency may exceed the duration of the audio itself, resulting in incoherent feedback audio, we use buffers to store input and output data and perform parallel signal processing in another independent thread. In this way, even if the processing delay is high due to algorithm complexity, the client can still hear coherent enhanced audio after a specific delay, thereby having a satisfactory service experience.
Figure 2: A schematic describes the experienced latency of this service.
Figure 2 describes the latency of the processing of the COG-MHEAR service. As shown in the system model, the client continuously collects voice and video information and sends it to the server. Assume that the length of the audio-video chunk sent by the client each time is $t_{\text{chunk}}$, that is, the client will send out audio and video after collecting a chunk of $t_{\text{chunk}}$. After receiving the data, the server will first perform preprocessing, such as feature extraction of video frames, and then store the processed data in the buffer. Since our algorithm tracks lips to help audio enhancement, the input of the algorithm must have a specific length $t_{i}$ to ensure sufficient visual features. When there are more than $t_{i}$ data in the buffer, the server will start executing the AVSE algorithm. Then, the server processes the incoming audio and video data sequentially, and we must set a sufficient buffer interval $t_{\Delta}$ to ensure that it can complete each process, which is given as
$t_{\Delta}>t_{a}$ (5)
where $t_{a}$ is the algorithm processing latency. The enhanced audio is stored in the output buffer, and the server begins feeding the enhanced audio back to the client when there is enough data in the output buffer. It should be noted that since $t_{i}$ is usually much larger than $t_{\Delta}$, if we want the initial audio to be enhanced by this algorithm, there will be an inevitable $t_{i}$-length waiting time for data, causing a high delay. So we play a trick, that is, we do not process the audio at the beginning of $t_{i}-t_{\Delta}$, thereby reducing the overall delay to
$t_{\text{delay}}=t_{1}+t_{2}+t_{3}=t_{\text{comm}}+t_{\text{proc}}$ (6)
where $t_{1}$ denotes the forward network latency, i.e., the delay incurred in transmitting data from the client to the server; $t_{2}$ denotes the server’s processing latency, i.e., the delay incurred by the server in processing the data; $t_{3}$ denotes the reverse network latency, i.e., the delay incurred in transmitting data from the server to the client. Thus, $t_{1}+t_{3}$ comprise the communication latency, and the processing latency $t_{\text{proc}}$ (i.e., $t_{2}$) can be decomposed into the buffer interval $t_{\Delta}$ and algorithm processing time $t_{a}$. As a result, the client can get coherent feedback audio $t_{\text{proc}}$ seconds after enabling the service, and audio enhancement is enabled after $t_{i}-t_{\Delta}$ seconds.
V Experimental Results
In this experiment, the auditory information undergoes a transformation via convolutional neural networks (CNNs), which can distill the audio patterns. The visual information is processed through the OpenCV library, which is an open-source computer vision and machine learning software library, to extract facial features from the face embeddings. It is implemented using the pre-trained Haar feature-based cascade classifiers, which are underpinned also by CNNs. Subsequently, the algorithm employs a fusion strategy where the extracted auditory and facial features coalesce, serving as inputs to a sophisticated long shortterm memory (LSTM) network. This step is crucial for integrating temporal dynamics and ensuring a coherent feature representation. The convergence of these modalities culminates in a series of FC layers, which tunes the composite features to reconstruct the target outputs. This framework provides a foundational blueprint, while the explicit architectural configuration (i.e., layers, neurons, etc.) can vary depending on the requirements and data pre-processing.
It should be noted that although we do buffering to ensure coherent audio feedback from the server, the capacity of the communication network is also extremely important to ensure reliable quality of service. The communication network needs to have sufficient capacity so that the round-trip delay of data transmission between the client and server is shorter than the time of the data itself to ensure a consistent service experience, denoted as
$t_{\text{comm}}\leq t_{\text{chunk}}$ (7)
Otherwise, after playing a chunk, if the data of the next chunk has not been fed back, there will be a blank period.
We measured the latency of a variety of networks in real-world environments. We record audio at a sampling rate of $16$ kHz and captured video frames at a resolution of $380\times 640$ pixels at a frame rate of $25$ fps. We set $t_{\text{chunk}}=40$ ms, congruent with the temporal extent of a single video frame. This configuration resulted in a raw data payload of approximately $0.3$ MB per audio-video chunk. Figure 3 shows the latency profiles observed across a suite of $100$ round-trip data transmissions. In our experimental setup, a desktop-based server and a laptop-based client were utilized to simulate a realistic user scenario. The communication between server and client was facilitated through Wi-Fi (802.11n, i.e., Wi-Fi 4), 4G, 5G, and a direct Ethernet cable link, respectively. We also tried the Amazon Web Services, which hosts the server codes, and the laptop-based client accesses it through the same Wi-Fi. The results suggest that only 5G network and direct Ethernet cable link could support a coherent service experience, i.e., satisfying $t_{\text{comm}}\leq t_{\text{chunk}}$ where $t_{\text{chunk}}=40$ ms in our experiments. Note that 4G and 5G network capacities may differ in different regions and with different carriers while choosing different AWS locations may also result in different latencies. We used Wi-Fi from the King’s Buildings, The University of Edinburgh (Edinburgh, UK), and for 4G and 5G network testing used a private network of Cambridge Wireless (Cambridge, UK). Using new Wi-Fi protocols, such as 802.11ac (Wi-Fi 5) or 802.11ax (Wi-Fi 6), can dramatically reduce the latency. In areas with poor signal coverage, we can compress the raw audio-video data to reduce the transmission load, obtaining reliable and consistent quality of service. Figure 4 shows the data size versus compression factors. It can be seen that proper compression of data can greatly reduce the data size and the required data rate, especially in the range of 80-100 compression factors. At a compression factor of 80, the data can be reduced to $\tilde{7}0$ KB with a trivial effect on the quality of the video frames. We also used a compression factor of 80 in our real-world demonstration in Cambridge to ensure smooth service, and experiments proved that the impact of other speech enhancement properties could be ignored and we could still achieve amazing noise reduction.
Figure 3: Network latency measured in real-world environments (round-trip latency for transferring $\tilde{0}.3$ MB of data) .
Figure 4: Audio-video data size versus compression factors.
V-A Algorithm Processing Latency
The communication latency has little effect on user experience delay ($t_{\text{delay}}$) but is critical to the coherence of feedback audio. And the real-world measurements show the ability to meet the demand using either the latest 5G or WI-Fi 6. In order to reduce the user experience latency, the most critical thing is to reduce the processing time of the algorithm, which can be achieved by reducing the input chunk size and the model size. However, since our algorithm needs to track lip vibrations in order to extract the target speaker’s voice from the mixed noise, decreasing the size of the input chunk usually incurs a performance loss. Figure 5 shows the algorithm processing latency versus the input chunk size in terms of frames, where one frame is $40$ ms. It can be seen that the processing delay is almost proportional to the input size. Theoretically, if we process each frame individually, then the processing latency will be less
Figure 5: Algorithm processing latency versus input chunk size.
than the duration of a frame, thus reducing the user experience latency to be within $100$ ms. However, reducing the input chunk size will result in a serious performance loss. Figure 6 plots the time-domain waveforms and spectrogram of both the noisy input audio and the output audio post-enhancement. The visualization is stratified by varying lengths of input chunks, clearly showcasing the interplay between chunk size and signal integrity. Speech enhanced with $10$s audio and video inputs yields a distinct and unblemished vocal output from the target speaker. When the enhancement is applied on a frame-by-frame basis, the resulting output is predominantly tainted with noise. A notable improvement in audio clarity is observed as we extend the duration of the input; however, this does not uniformly guarantee optimal performance across all temporal segments. Some intervals exhibit satisfactory enhancement, while others remain suboptimal. Consequently, to achieve a consistently high Quality of Service, we found it necessary to utilize 10-second blocks of audio and video for each processing iteration, which resulted in about $2.2$s of algorithm processing time, and to make sure that the server could complete each processing, we set an interval of 2.32.42.3-2.4s, which resulted in a total latency of about $5$s.
Figure 6: Time plot and spectrogram of the input noisy audio and output enhanced audio processed with different input chunk lengths.
V-B Algorithm development
To mitigate the inherent latency in our system, we tried to refine our network’s proficiency in distilling potent feature representations from short segments of audio and visual data. To this end, we have calibrated the network training scheme to work with condensed input durations of merely $200$ ms. Leveraging the architecture depicted in LABEL:figure1, we constructed a series of networks, reducing their structures by paring down layers and neural nodes with the express aim of shortening processing delays. Table I lists the total parameters and processing latency of each of the three networks.
Models Total parameters Time complexity
Model 1 1,540,396 (5.88 MB) $\tilde{1}200$ ms
Model 2 603,564 (2.30 MB) $\tilde{5}50$ ms
Model 3 202,564 (791.27 KB) $\tilde{3}50$ ms
TABLE I: MODELS’ PARAMETERS AND COMPLEXITY
Figure 7 plots the time-domain waveforms and spectrogram of the original target speaker’s voice, noisy input audio, and the output audio post-enhancement. The reduction in visual information due to the shortened input duration results in a lower reproduction performance of the original sound when it is weak, but the noise reduction is still satisfactory.
Figure 7: Time plot and spectrogram of the input noisy audio and output enhanced audio processed with different network architectures.
VI CONCLUSION
In conclusion, this paper’s exploration of an Audio-Video-Based Speech Enhancement (AVSE) system underscores its potential to revolutionize speech recognition by integrating audio and visual data through advanced deep learning techniques. Among the myriad of practical applications, one of the most impactful is in the development of next-generation hearing aids. These advanced hearing aids, equipped with AVSE capabilities, could offer a transformative experience for individuals with hearing impairments, providing them with an enhanced ability to understand speech in noisy environments by leveraging both auditory and visual cues.
Looking to the future, several promising research directions could further elevate AVSE technology. Optimizing the balance between audio and visual modalities is crucial, as it can vary depending on environmental factors and the specific needs of the application. Investigating diverse DNN architectures can potentially unveil more effective ways of processing and merging audio-visual data. Moreover, experimenting with various fusion functions beyond simple concatenation could lead to more nuanced and context-aware integration of sensory inputs.
The broader implications of AVSE technology extend to a variety of real-world scenarios. Its application in military communications can drastically improve operational efficiency on the battlefield. In construction sites or any noisy industrial environment, AVSE can facilitate clearer internal communications among workers. The technology also has significant applications in videoconferencing, journalism, and sports, particularly in enhancing communication in noisy stadiums.
Each application presents unique challenges, suggesting a need for AVSE solutions that are customized to specific environments and user requirements. Future advancements in this field not only promise to push the boundaries of speech enhancement technologies but also offer substantial improvements in diverse sectors that rely on clear and effective communication. The integration of AVSE systems into hearing aids represents a particularly profound opportunity to enrich the lives of individuals with hearing impairments, marking a significant step forward in assistive technology.
Acknowledgements