### v2

# Audio–Visual Speech Enhancement: Architectural Design and Deployment Strategies

Abstract
This paper presents a real-time wireless 5G cloud-based audio-visual speech enhancement (AVSE) framework that integrates convolutional neural networks (CNNs) for acoustic feature extraction and OpenCV-based facial embedding for visual processing. The extracted multimodal features are fused through a long short-term memory (LSTM) network to reconstruct clean speech from noisy input. Experimental evaluations across multiple network configurations—Wi-Fi, 4G, 5G, Ethernet, and AWS cloud—demonstrate that only 5G and direct Ethernet links consistently satisfy the coherence condition ($t_{\text{comm}}\leq t_{\text{chunk}}$) for continuous streaming. By applying an 80× compression factor, the payload per 40 ms audio-video chunk is reduced from approximately 0.3 MB to 70 KB with minimal quality loss, enabling reliable transmission under bandwidth constraints. Results further reveal a trade-off between algorithmic latency and enhancement performance: smaller models and shorter input windows lower processing time but reduce speech clarity. The proposed architecture achieves end-to-end latency of approximately 5 s with strong noise reduction and speech intelligibility, demonstrating the feasibility of deploying multimodal speech enhancement in cloud-assisted, low-latency wireless environments.
Index Terms:
Audio-Visual Speech Enhancement, 5G Networks, Cloud Computing, Deep Learning, CNN-LSTM Fusion, Real-Time Processing, Wireless Communication, Multimodal Signal Processing
I Introduction
In many real-world acoustic environments, speech signals are severely degraded by additive background noise, reverberation, and competing speakers, which substantially reduce intelligibility and perceived quality. Conventional audio-only speech enhancement techniques aim to recover a clean speech signal ($x[n]$) from a noisy observation ($y[n]=x[n]+d[n]$), where ($d[n]$) denotes noise or interference. However, when the signal-to-noise ratio (SNR) is low or overlapping talkers are present, these audio-only approaches often struggle to reliably recover intelligible speech.
Human listeners naturally exploit not only auditory cues but also visual information — such as lip movements and facial articulations — to compensate for challenging acoustic conditions. This perceptual phenomenon is well known (e.g., the McGurk effect) and underscores the value of multimodal input in speech perception. Visual speech cues remain unaffected by acoustic masking and therefore offer complementary information that is particularly valuable when the acoustic channel is corrupted.
Audio-visual speech enhancement (AVSE) builds on this insight by jointly leveraging both audio and visual modalities (for example, microphone audio plus camera capture of the speaker’s lips) to enhance or recover target speech. Deep-learning based AVSE systems typically extract acoustic features (e.g., log-mel spectrograms, STFT magnitudes) and visual features (e.g., lip region embeddings via CNNs), and then fuse them in a neural network to generate enhanced speech. Surveys (e.g., Michelsanti et al. 2021 [michelsanti2021overview]) have shown that multimodal AVSE systems often outperform audio-only baselines, especially in low-SNR or multi-speaker conditions.
Despite these algorithmic and modelling advances, deploying AVSE in real-world applications introduces a distinct set of system-level challenges. Wearable devices such as hearing aids or smart glasses must operate under tight constraints of latency, power consumption, and wireless connectivity. They also need to handle streaming audio and video data in real time. Offloading heavy deep‐learning inference to a remote cloud or edge server emerges as a feasible solution when the local device is resource‐constrained.
In this context, the advent of next-generation wireless networks (e.g., 5G NR) with high bandwidth and ultra-low latency offers a promising platform for real-time AVSE deployment. For example, recent demonstrations of 5G/IoT enabled AVSE for multimodal hearing aids illustrate the viability of high-data-rate, low-latency uplinks combined with cloud-based enhancement [gupta20235g]. By harnessing the 5G link and off-loading computation to the cloud, devices can remain light, efficient and power-conserving, while still benefiting from powerful multimodal models.
However, the full stack of implementation trade‐offs—including uplink/downlink latency, wireless reliability under mobility, compute placement (device vs edge vs cloud), streaming bandwidth for visual + audio, and real-time intelligibility/quality outcomes—remains under-explored in AVSE research. The present work addresses this gap by proposing and implementing a wireless 5G-based cloud architecture for AVSE. Our system captures lip-region video and noisy speech on a wearable front‐end, transmits the data via a 5G uplink to a remote cloud server, performs multimodal deep-learning AVSE processing, and streams back the enhanced speech to the device. We evaluate the system in terms of end-to-end latency, wireless throughput, model inference time, and speech intelligibility/quality under realistic noisy and multi-speaker conditions. In so doing, we extend the state of the art from model-centric AVSE research to a full-stack, deployment-oriented system suitable for mobile and wearable real-time use.
The remainder of this paper is structured as follows. Section II reviews related work on AVSE, wireless/5G off‐load implementations and deployment strategies. Section III presents the system design and implementation of our 5G cloud-based AVSE pipeline. Section IV reports experimental evaluation results and discusses findings and practical implications. Finally, Section V concludes the paper and outlines future research directions.
II Related Work
Research in audio-visual speech enhancement (AVSE) has progressed substantially over the past decade, particularly with the proliferation of deep learning and multimodal approaches. Early work in audio-visual speech separation, such as the seminal model by Ephrat et al. (“Looking to Listen at the Cocktail Party”), demonstrated that a joint audio‐visual network significantly outperforms audio‐only baselines in multi‐speaker and noisy conditions by leveraging lip‐motion cues [ephrat2018looking]. Subsequently, surveys such as Michelsanti et al. (2021) [michelsanti2021overview] provided a systematic overview of the field, highlighting key components such as acoustic feature extraction, visual feature encoding, fusion strategies (early vs intermediate vs late), and standard objective functions for AVSE/AVSS.
In more algorithm‐centric developments, attention has turned to how to fuse modalities effectively and design architectures that can operate in real‐time or near‐real‐time. For example, Xu and Hao [xu2021multi] proposed the attentional multi‐layer fusion convolution network (AMFFCN) to apply soft‐threshold attention at each layer of the decoder to select informative modalities. More recently, Gao and Grauman introduced the VisualVoice framework, which learns cross-modal embeddings (face appearance plus audio) for improved speech separation in unconstrained video scenes [gao2021visualvoice].
In terms of system implementation and deployment, particularly for wearable or real‐world applications, researchers have begun to tackle not just the algorithmic enhancement but also the architectural trade-offs of where the processing is done (device, edge, or cloud). Adeel et al. (2018) [adeel2018real] envisioned a 5G IoT enabled lip-reading driven hearing‐aid framework that off-loads AV data to the cloud, employing lightweight encryption and aiming for round‐trip latency targets on the order of a few milliseconds. On the dataset and baseline front, Michelsanti’s survey [michelsanti2021overview] underscores the importance of dataset design and evaluation metrics (e.g., STOI, PESQ, SI-SDR) when comparing AVSE systems.
Despite algorithmic advances, few works systematically compare deployment strategies (cloud vs edge vs device) under realistic network and computing constraints. To fill this gap, our work develops a comprehensive AVSE architecture (CNN + LSTM fusion) and performs a comparative analysis across deployment scenarios under Ethernet, Wi-Fi, 4G and 5G conditions. Although, prior work has established the foundational value of adding visual cues to speech enhancement; explored a wide variety of fusion architectures and benchmarks; and begun to address deployment challenges, The full spectrum of practical trade‐offs—end‐to‐end latency, network delay, compute location, and real‐world system performance—remains less explored, motivating the present implementation and experimental‐study contribution.
III Proposed Solution
Figure 1 presents a pragmatic framework for the deployment of the developed AVSE algorithm within the context of real-world applications. In order to enable the service in low quality terminals (e.g., IoT devices), the remote server takes over all the processing, including the pre-processing of the raw data. By doing so, any device with a camera, microphone, and communication capabilities can use the service without any requirements on its computing power. This allows us to apply the service to some cheap IoT devices to achieve specific functions. For high-performance devices (e.g., smartphones), offloading all processing to the server can also reduce energy consumption and improve battery life.
Figure 1: A block diagram of enabling real-world COG-MHEAR service on terminal devices.
As the figure depicts, the client terminals only collect the raw audio and video, then upload them to the server and wait for the feedback from the server. It should be noted that collecting and playing audio requires parallel processing, so we use two independent threads. The server will listen to clients and wait for the raw audio-video data from clients. It starts processing using the stored trained model once it listens to data transmission from the client. Considering that processing latency may exceed the duration of the audio itself, resulting in incoherent feedback audio, we use buffers to store input and output data and perform parallel signal processing in another independent thread. In this way, even if the processing delay is high due to algorithm complexity, the client can still hear coherent enhanced audio after a specific delay, thereby having a satisfactory service experience.
Figure 2: A schematic describes the experienced latency of this service.
Figure 2 describes the latency of the processing of the COG-MHEAR service. As shown in the system model, the client continuously collects voice and video information and sends it to the server. Assume that the length of the audio-video chunk sent by the client each time is $t_{\text{chunk}}$, that is, the client will send out audio and video after collecting a chunk of $t_{\text{chunk}}$. After receiving the data, the server will first perform preprocessing, such as feature extraction of video frames, and then store the processed data in the buffer. Since our algorithm tracks lips to help audio enhancement, the input of the algorithm must have a specific length $t_{i}$ to ensure sufficient visual features. When there are more than $t_{i}$ data in the buffer, the server will start executing the AVSE algorithm. Then, the server processes the incoming audio and video data sequentially, and we must set a sufficient buffer interval $t_{\Delta}$ to ensure that it can complete each process, which is given as
$t_{\Delta}>t_{a}$ (1)
where $t_{a}$ is the algorithm processing latency. The enhanced audio is stored in the output buffer, and the server begins feeding the enhanced audio back to the client when there is enough data in the output buffer. It should be noted that since $t_{i}$ is usually much larger than $t_{\Delta}$, if we want the initial audio to be enhanced by this algorithm, there will be an inevitable $t_{i}$-length waiting time for data, causing a high delay. So we play a trick, that is, we do not process the audio at the beginning of $t_{i}-t_{\Delta}$, thereby reducing the overall delay to
$t_{\text{delay}}=t_{1}+t_{2}+t_{3}=t_{\text{comm}}+t_{\text{proc}}$ (2)
where $t_{1}$ denotes the forward network latency, i.e., the delay incurred in transmitting data from the client to the server; $t_{2}$ denotes the server’s processing latency, i.e., the delay incurred by the server in processing the data; $t_{3}$ denotes the reverse network latency, i.e., the delay incurred in transmitting data from the server to the client. Thus, $t_{1}+t_{3}$ comprise the communication latency, and the processing latency $t_{\text{proc}}$ (i.e., $t_{2}$) can be decomposed into the buffer interval $t_{\Delta}$ and algorithm processing time $t_{a}$. As a result, the client can get coherent feedback audio $t_{\text{proc}}$ seconds after enabling the service, and audio enhancement is enabled after $t_{i}-t_{\Delta}$ seconds.
IV Experimental Results
In this experiment, the auditory information undergoes a transformation via convolutional neural networks (CNNs), which can distill the audio patterns. The visual information is processed through the OpenCV library, which is an open-source computer vision and machine learning software library, to extract facial features from the face embeddings. It is implemented using the pre-trained Haar feature-based cascade classifiers, which are underpinned also by CNNs. Subsequently, the algorithm employs a fusion strategy where the extracted auditory and facial features coalesce, serving as inputs to a sophisticated long shortterm memory (LSTM) network. This step is crucial for integrating temporal dynamics and ensuring a coherent feature representation. The convergence of these modalities culminates in a series of FC layers, which tunes the composite features to reconstruct the target outputs. This framework provides a foundational blueprint, while the explicit architectural configuration (i.e., layers, neurons, etc.) can vary depending on the requirements and data pre-processing.
It should be noted that although we do buffering to ensure coherent audio feedback from the server, the capacity of the communication network is also extremely important to ensure reliable quality of service. The communication network needs to have sufficient capacity so that the round-trip delay of data transmission between the client and server is shorter than the time of the data itself to ensure a consistent service experience, denoted as
$t_{\text{comm}}\leq t_{\text{chunk}}$ (3)
Otherwise, after playing a chunk, if the data of the next chunk has not been fed back, there will be a blank period.
We measured the latency of a variety of networks in real-world environments. We record audio at a sampling rate of $16$ kHz and captured video frames at a resolution of $380\times 640$ pixels at a frame rate of $25$ fps. We set $t_{\text{chunk}}=40$ ms, congruent with the temporal extent of a single video frame. This configuration resulted in a raw data payload of approximately $0.3$ MB per audio-video chunk. Figure 3 shows the latency profiles observed across a suite of $100$ round-trip data transmissions. In our experimental setup, a desktop-based server and a laptop-based client were utilized to simulate a realistic user scenario. The communication between server and client was facilitated through Wi-Fi (802.11n, i.e., Wi-Fi 4), 4G, 5G, and a direct Ethernet cable link, respectively. We also tried the Amazon Web Services, which hosts the server codes, and the laptop-based client accesses it through the same Wi-Fi. The results suggest that only 5G network and direct Ethernet cable link could support a coherent service experience, i.e., satisfying $t_{\text{comm}}\leq t_{\text{chunk}}$ where $t_{\text{chunk}}=40$ ms in our experiments. Note that 4G and 5G network capacities may differ in different regions and with different carriers while choosing different AWS locations may also result in different latencies. We used Wi-Fi from the King’s Buildings, The University of Edinburgh (Edinburgh, UK), and for 4G and 5G network testing used a private network of Cambridge Wireless (Cambridge, UK). Using new Wi-Fi protocols, such as 802.11ac (Wi-Fi 5) or 802.11ax (Wi-Fi 6), can dramatically reduce the latency. In areas with poor signal coverage, we can compress the raw audio-video data to reduce the transmission load, obtaining reliable and consistent quality of service. Figure 4 shows the data size versus compression factors. It can be seen that proper compression of data can greatly reduce the data size and the required data rate, especially in the range of 80-100 compression factors. At a compression factor of 80, the data can be reduced to $\tilde{7}0$ KB with a trivial effect on the quality of the video frames. We also used a compression factor of 80 in our real-world demonstration in Cambridge to ensure smooth service, and experiments proved that the impact of other speech enhancement properties could be ignored and we could still achieve amazing noise reduction.
Figure 3: Network latency measured in real-world environments (round-trip latency for transferring $\tilde{0}.3$ MB of data) .
Figure 4: Audio-video data size versus compression factors.
IV-A Algorithm Processing Latency
The communication latency has little effect on user experience delay ($t_{\text{delay}}$) but is critical to the coherence of feedback audio. And the real-world measurements show the ability to meet the demand using either the latest 5G or WI-Fi 6. In order to reduce the user experience latency, the most critical thing is to reduce the processing time of the algorithm, which can be achieved by reducing the input chunk size and the model size. However, since our algorithm needs to track lip vibrations in order to extract the target speaker’s voice from the mixed noise, decreasing the size of the input chunk usually incurs a performance loss. Figure 5 shows the algorithm processing latency versus the input chunk size in terms of frames, where one frame is $40$ ms. It can be seen that the processing delay is almost proportional to the input size. Theoretically, if we process each frame individually, then the processing latency will be less
Figure 5: Algorithm processing latency versus input chunk size.
than the duration of a frame, thus reducing the user experience latency to be within $100$ ms. However, reducing the input chunk size will result in a serious performance loss. Figure 6 plots the time-domain waveforms and spectrogram of both the noisy input audio and the output audio post-enhancement. The visualization is stratified by varying lengths of input chunks, clearly showcasing the interplay between chunk size and signal integrity. Speech enhanced with $10$s audio and video inputs yields a distinct and unblemished vocal output from the target speaker. When the enhancement is applied on a frame-by-frame basis, the resulting output is predominantly tainted with noise. A notable improvement in audio clarity is observed as we extend the duration of the input; however, this does not uniformly guarantee optimal performance across all temporal segments. Some intervals exhibit satisfactory enhancement, while others remain suboptimal. Consequently, to achieve a consistently high Quality of Service, we found it necessary to utilize 10-second blocks of audio and video for each processing iteration, which resulted in about $2.2$s of algorithm processing time, and to make sure that the server could complete each processing, we set an interval of 2.32.42.3-2.4s, which resulted in a total latency of about $5$s.
Figure 6: Time plot and spectrogram of the input noisy audio and output enhanced audio processed with different input chunk lengths.
IV-B Algorithm development
To mitigate the inherent latency in our system, we tried to refine our network’s proficiency in distilling potent feature representations from short segments of audio and visual data. To this end, we have calibrated the network training scheme to work with condensed input durations of merely $200$ ms. Leveraging the architecture depicted in Figure 7, we constructed a series of networks, reducing their structures by paring down layers and neural nodes with the express aim of shortening processing delays. Table I lists the total parameters and processing latency of each of the three networks.
Figure 7: Audio-Visual Speech Enhancement AI model architecture.
Models Total parameters Time complexity
Model 1 1,540,396 (5.88 MB) $\tilde{1}200$ ms
Model 2 603,564 (2.30 MB) $\tilde{5}50$ ms
Model 3 202,564 (791.27 KB) $\tilde{3}50$ ms
TABLE I: MODELS’ PARAMETERS AND COMPLEXITY
Figure 8 plots the time-domain waveforms and spectrogram of the original target speaker’s voice, noisy input audio, and the output audio post-enhancement. The reduction in visual information due to the shortened input duration results in a lower reproduction performance of the original sound when it is weak, but the noise reduction is still satisfactory.
Figure 8: Time plot and spectrogram of the input noisy audio and output enhanced audio processed with different network architectures.
V CONCLUSION
This study demonstrated a complete 5G cloud-based audio-visual speech enhancement (AVSE) system that integrates CNN-based acoustic processing and OpenCV facial feature extraction with an LSTM fusion network for temporal coherence. Experimental results showed that real-time performance depends critically on network latency and algorithmic processing time. Only 5G and wired Ethernet consistently met the coherence requirement ($t_{\text{comm}}\leq t_{\text{chunk}}$) (40 ms) for uncompressed 0.3 MB audio-video chunks, while compression at a factor of 80 reduced payloads to about 70 KB with negligible degradation, enabling smoother operation in constrained networks.
The findings highlight that network quality, chunk size, and model complexity jointly determine system responsiveness and enhancement quality. Although shorter input windows and smaller models reduce processing latency—from about 1.2 s to 0.35 s—they also degrade speech reconstruction, particularly under weak target signals. Achieving optimal performance therefore requires balancing temporal context and computational efficiency. Future work will focus on adaptive chunking, lightweight streaming architectures, and edge-assisted processing to further minimize delay while preserving enhancement fidelity in real-world wireless environments.
Acknowledgements
This work was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) Grant No. EP/XXXXXXX/X. The authors would like to gratefully acknowledge Haifeng Luo from the University of XXXX for his contribution in implementing XXXX.